OpenStudy (anonymous):

could you please help me with three math problem 1.lx-4yl x=-3 and y=2/4, 2. 3xsquarey-6xysquare-9xy and 3. 5msquare+5nsquare.

OpenStudy (netlopes1):

Is this your question? \[1.\quad |x-4y|, \,\,x=-3\,\, and\,\, y=\frac{1}{2}.\\ 2.\quad 3x^{2}y-6xy^{2}-9xy.\\ 3.\quad 5m^{2}+5n^{2}. \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=2/4

OpenStudy (netlopes1):

1. |-3-4.(2/4)| or |-3-2| = |-5| = 5 (final)

OpenStudy (netlopes1):

2. May I use the same values x=-3 and y=2\4?

OpenStudy (anonymous):

no two and three is factoring each problem

OpenStudy (netlopes1):

ok...

OpenStudy (netlopes1):

\[2.\quad3x^{2}y-6xy^{2}-9xy\rightarrow 3xy(x-2y-3)\quad (final)\]

OpenStudy (netlopes1):

\[3.\quad 5m^{2}+5n^{2}\rightarrow 5(m^{2}+n^{2} (final)\]

OpenStudy (netlopes1):

\[3.\quad 5m^{2}+5n^{2}\rightarrow 5(m^{2}+n^{2}) (better)\]

OpenStudy (netlopes1):

Thank you for your trust!!

OpenStudy (anonymous):

could you help me with this and show me the work 4c+4d-as-ad=

OpenStudy (netlopes1):

ok.... 4(c+d)-a(s+d) (final)